calculate ph of 0.200 mol/l of H3PO4 (aq) if ka = 2.00x10-2


1
Expert's answer
2021-02-17T07:47:17-0500

The reaction is:-


"H_3PO_4+H_2O\rightarrow H_3O^++H_2PO_4^-"

"\text{at t }=0" "0.2" 0 0


Let at time t' The concentration of "H_3O^+" became x


"\text{ at t }=t'" "0.2-x" "x" "x"


So, "k_a=\dfrac{[H_3O^+][H_2PO_4^-]}{[H_3PO_4]}"


"2\times 10^{-2}=\dfrac{x.x}{0.2-x}"


Solving for "x:"


"0.02(0.2-x)=x^2"


"\Rightarrow x^2+0.02x-0.004=0"


From above we get "x=0.01"


"[H_3O^+]=0.01"


"P_H=-log[H_3O^+]"


"=-log(0.01)

 \\=2"


Hence "P_H" is 2.